There are 3 football teams the first team has 10 points more than the second team and the second team has 5 more than the last team all together they have 110 points
Answer by stanbon(75887)

There are 3 football teams the first team has 10 points more than the second team and the second team has 5 more than the last team all together they have 110 points
----
Equations:
a = b+10
b = c+5
a + b + c = 110
----------------------------
Substitute and solve for "a"::
a + (a-10) + (b-5) = 110
a + (a-10) + (a-10-5) = 110
-----
3a -25 = 110
3a = 135
a = 45
b = a-10 = 35
c = b-5 = 30
-----------------
Cheers,
Stan H.
-------------